# Phase domain walls in coherently driven Bose-Einstein condensates

### v3

@@ V.2 Spontaneous phenomena
Figure 10: Typical stages of evolution of a system with zero initial conditions under a very slow increase in $f$. (a) Highly disordered state soon after the spin symmetry breakdown. (b) Onset of phase domains. (c) Self-organization of nonstationary domains. (d) Eventual static solution at $f^{2}/f_{2}^{2}\approx 0.14$. The color map represents $\phi_{+}$. Parameters: $D=\Omega$, $\gamma/\Omega=5\times 10^{-3}$, $R=125\,\xi$.
@@
When $f$ is increased further, the system exhibits dynamically stable domain walls with $q=0$. Notice that even Fig. 3 suggests that the $q=0$ states become less steep and, thus, less energetically costly than the $q=\pm 1$ ones as $f$ increases. Figure 11 shows a crossover solution in which the two types of domain walls coexist. It evolved from zero initial conditions, similarly to Fig. 10, under increasing $f^{2}$ up to  ${\sim}\,0.16\,f_{2}^{2}$. The $q=0$ part of the domain wall is seen as a straight line segment characterized by opposite $\phi_{+}$ and $\phi_{-}$ [Figs. 11(a) and 11(b)], high $S_{1}$ [Fig. 11(c)], and almost zero $S_{3}$ [Fig. 11(d)]. The rectilinear shape is indicative of the domain-wall tension that has an energetic origin in analogy to the $2\pi$ domain walls in atomic BECs [3, 7].
@@ Appendix A Parameters of the polariton system
The set of parameters used in our numerical examples corresponds to a GaAs-based microcavity: $\epsilon=12$, $E_{0}=1.5$ eV, $\mathcal{R}=10$ meV, $D=\Omega=1$ meV; $\gamma$ is small and chosen arbitrarily. The only unconventional assumption is that $\Omega$ largely exceeds $\gamma$. Calculations show that phase domains spontaneously form in some range of $f$ at $\Omega/\gamma\gtrsim 10$ (see also a discussion in [50]). However, in order to robustly observe both types of domain walls, one has to further increase $\Omega/\gamma$ up to  ${\sim}\,10^{2}$; for instance, Figs. 6, 10, and 11 are obtained at $\Omega/\gamma=200$. In practice, high $\Omega/\gamma$ could originate from a structural anisotropy (e. g., owing to an in-plane mechanical stress [20] or magnetic field) combined with a high quality factor $Q\approx E_{0}/2\gamma$. These conditions seem to be fairly achievable in the immediate future.
@@ Appendix B All vacuum states and elementary excitations
Once this threshold is reached, $u$ starts growing spontaneously (even if $f$ is kept constant) up to the magnitude $u\gtrsim 2D-\Omega$ that corresponds the alternative root of (28) for the same $f$. Such a switch between two vacuum states, mediated by transient excitation of many modes with $\mathbf{k}\neq 0$, was studied earlier in the case of a scalar bistable system [49]. The two-component system with $D\sim\Omega\gg\gamma$ behaves differently in that its upper stationary solution obeying (28) is unstable with respect to the spin symmetry breaking. Technically, one can make sure that the eigenstates $\zeta_{\pm}$ and $\eta_{\pm}$ corresponding to $\mathop{\mathrm{Im}}\tilde{E}>0$ are antisymmetric ($\zeta_{+}=-\zeta_{-}$, $\eta_{+}=-\eta_{-}$), so their spontaneous amplification necessarily breaks the spin symmetry (see Sec. V.2).
@@
If $D\sim\Omega\gg\gamma$, quartic equation $df/du=0$ has four real-valued roots, of which two, $r_{1,2}$, fall within $(l_{1},l_{2})$. Let $r_{1}<r_{2}$, then the first relevant interval $l_{1}<u<r_{1}$ has a positive slope $(du/df>0)$ and contains the spin doublets whose stability (sign of $\Gamma$) depends on $f$. The next one, $r_{1}<u<r_{2}$, is wholly unstable ($du/df<0$ and $\Gamma>0$). The last interval $r_{2}<u<l_{2}$ contains the opposite-phase doublets which, by contrast, are stable for each $f$. At $u\to l_{2}$, both $u_{+}-u_{-}$ and $\alpha_{+}-\alpha_{-}$ tend to zero, (33) coincides with (28), and (37) coincides with (29); moreover, exactly at this point the spin-symmetric solutions become stable again, displaying the transcritical bifurcation.
@@
It is seen that $f_{2}\gg f_{1}$ at $\gamma\to 0$, however, the overall change in $u$ throughout the top branch is very small $({\sim}\,\gamma)$. Strictly speaking, (33) is senseless at $\gamma=0$ because of the $0/0$ indeterminacy in its last term, which means that $u$ becomes independent of $f$. However, everything remains regular at $\gamma\to 0^{+}$. Below, we always suppose that $D\sim\Omega\gg\gamma$, thus, we can substitute $u=2\mu-\rho$ directly into (33) and then drop the $\rho^{4}$ and $\rho^{3}$ terms, which eventually yields
@@
It turns out that $u_{+}$ and $u_{-}$ are very close almost irrespective of $f$: even the maximum $|S_{3}|$ reached at $f=f_{1}$ is of the order of $\sqrt{\gamma/\mu}$. On the other hand, it is clear from (36) that condition $u_{+}=u_{-}$ necessarily implies $e^{i\alpha_{+}}=e^{-i\alpha_{-}}$ as long as $\gamma\to 0$. As a result, from (35) and (36) we have
@@
for two solutions in the phase doublet (generally, the sign of $\sin(\alpha_{-}-\alpha_{+})$ is equal to the sign of $S_{3}$). The fact that both components are $\pi$-shifted relative to the driving field at $f=f_{2}$ agrees with the spin-symmetric model in the range of great $f$ and $u$.

### v4

@@ V.2 Spontaneous phenomena
Figure 10: Typical stages of evolution of a system with zero initial conditions under a slow increase of $f$. (a) Highly disordered state soon after the spin symmetry breakdown. (b) Onset of phase domains. (c) Self-organization of nonstationary domains. (d) Eventual static solution at $f^{2}/f_{2}^{2}\approx 0.14$. The color map represents $\phi_{+}$. Parameters: $D=\Omega$, $\gamma/\Omega=5\times 10^{-3}$, $R=125\,\xi$. The dynamics is shown in Supplemental Material V [56].
@@
The system evolution outlined in Fig. 10 is explicitly shown in Supplemental Material V  [^5] See Supplemental Material V for a video presentation corresponding to Fig. 10.. In this calculation, the pump intensity $f^{2}(t)$ was increased linearly during the time period of $T\approx 450\,\hbar/\mu$ and then kept constant. To reduce the time of the spin symmetry breaking, we added to $f$ at $t<T$ a weak complex white noise $\tilde{f}_{\pm}(\mathbf{r},t)$ with a fixed amplitude $|\tilde{f}_{\pm}|=\beta$ and zero mean over both $\mathbf{r}$ and $t$. Since the symmetry breaks down owing to the exponential growth of excitations, the time of this process shows only a logarithmic increase with decreasing $\beta$. Thus, the obtained results are qualitatively independent of $T$ and $\beta$ (in particular, a disordered state as in Fig. 10(a) is inevitable).
@@
When $f$ is increased further compared to Fig. 10, the system exhibits dynamically stable domain walls with $q=0$. Notice that even Fig. 4 suggests that the $q=0$ states become less steep and, thus, less energetically costly than the $q=\pm 1$ ones as $f$ increases. Figure 11 shows a crossover solution in which the two types of domain walls coexist. It evolved from zero initial conditions, similarly to Fig. 10, under increasing $f^{2}$ up to  ${\sim}\,0.16\,f_{2}^{2}$. The $q=0$ part of the domain wall is seen as a straight line segment characterized by opposite $\phi_{+}$ and $\phi_{-}$ [Figs. 11(a) and 11(b)], high $S_{1}$ [Fig. 11(c)], and almost zero $S_{3}$ [Fig. 11(d)]. The rectilinear shape is indicative of the domain-wall tension that has an energetic origin in analogy to the $2\pi$ domain walls in atomic BECs [3, 6].
@@
Experimentally, phase domains and both types of domain walls can be distinguished by measuring the Stokes vector components of the emitted radiation. The domains per se are distinguished by the sign of $S_{2}\approx\pm\sin 2\alpha$ [Eq. (9)]; the $q=0$ domain walls have $S_{1}=+1$ and $S_{3}=0$; the q=1q=-1 ($q=+1$) domain walls have $S_{1}\approx-1$ and positive (negative) $S_{3}$. These signatures are clearly seen in Figs. 11(c) and (d); the signs of $S_{3}$ for the $q=\pm 1$ domain walls are in line with Fig. 3.
@@ Appendix A Parameters of the polariton system
The set of parameters used in our numerical examples corresponds to a GaAs-based microcavity: $\epsilon=12$, $E_{0}=1.5$ eV, $\mathcal{R}=10$ meV, $D=\Omega=1$ meV; $\gamma$ is small and chosen arbitrarily. The only unconventional assumption is that $\Omega$ largely exceeds $\gamma$. Calculations show that phase domains spontaneously form in some range of $f$ at $\Omega/\gamma\gtrsim 10$ (see also a discussion in [58]). However, in order to robustly observe both types of domain walls, one has to further increase $\Omega/\gamma$ up to  ${\sim}\,10^{2}$; for instance, Figs. 6, 10, and 11 are obtained at $\Omega/\gamma=200$. In practice, high $\Omega/\gamma$ could originate from a structural anisotropy (e. g., owing to an in-plane mechanical stress [23] or magnetic field) combined with a high quality factor $Q\approx E_{0}/2\gamma$. These conditions seem to be achievable in the immediate future.
@@ Appendix B All vacuum states and elementary excitations
Once this threshold is reached, $u$ starts growing spontaneously (even if $f$ is kept constant) up to the magnitude $u\gtrsim 2D-\Omega$ that corresponds to a different root of (26) for the same $f$. Such a switch between two vacuum states, mediated by transient excitation of many modes with $\mathbf{k}\neq 0$, was studied earlier in the case of a scalar bistable system [57]. The two-component system with $\gamma\ll D\sim\Omega$ behaves differently in that its upper stationary solution obeying (26) is unstable with respect to the spin symmetry breaking. Technically, one can make sure that the eigenstates $\zeta_{\pm}$ and $\eta_{\pm}$ corresponding to $\mathop{\mathrm{Im}}\tilde{E}>0$ are antisymmetric ($\zeta_{+}=-\zeta_{-}$, $\eta_{+}=-\eta_{-}$), so their spontaneous amplification necessarily breaks the spin symmetry (see Sec. V.2).
@@
If $\gamma\ll D\sim\Omega$, quartic equation $df/du=0$ has four real-valued roots, of which two, $r_{1,2}$, fall within $(l_{1},l_{2})$. Let $r_{1}<r_{2}$, then the first relevant interval $l_{1}<u<r_{1}$ has a positive slope $(du/df>0)$ and contains the spin doublets whose stability (sign of $\Gamma$) depends on $f$. The next one, $r_{1}<u<r_{2}$, is wholly unstable ($du/df<0$ and $\Gamma>0$). The last interval $r_{2}<u<l_{2}$ contains the opposite-phase doublets which, by contrast, are stable for each $f$. At $u\to l_{2}$, both $u_{+}-u_{-}$ and $\alpha_{+}-\alpha_{-}$ tend to zero, (31) coincides with (26), and (35) coincides with (27); moreover, exactly at this point the spin-symmetric solutions become stable again, displaying the transcritical bifurcation.
@@
It is seen that $f_{2}\gg f_{1}$ at $\gamma\to 0$, however, the overall change in $u$ throughout the top branch is very small $({\sim}\,\gamma)$. Strictly speaking, (31) is senseless at $\gamma=0$ because of the $0/0$ indeterminacy in its last term, which means that $u$ becomes independent of $f$. However, everything remains regular at $\gamma\to 0^{+}$. Below, we always suppose that $\gamma\ll D\sim\Omega$, thus, we can substitute $u=2\mu-\rho$ directly into (31) and then drop the $\rho^{4}$ and $\rho^{3}$ terms, which eventually yields
@@
It turns out that $u_{+}$ and $u_{-}$ are very close almost irrespective of $f$ (even the maximum $|S_{3}|$ reached at $f=f_{1}$ is of the order of $\sqrt{\gamma/\mu}$). On the other hand, it is clear from (34) that condition $u_{+}=u_{-}$ necessarily implies $e^{i\alpha_{+}}=e^{-i\alpha_{-}}$ as long as $\gamma\to 0$. As a result, from (33) and (34) we have
@@
for two solutions in the phase doublet (generally, the sign of $\sin(\alpha_{-}-\alpha_{+})$ is equal to the sign of $S_{3}$). The fact that both components are $\pi$-shifted relative to the driving field at $f=f_{2}$ agrees with the spin-symmetric model at great $f$ and $u$.